# Bayesian inferences on covariant density functionals from multimessenger astrophysical data: Nucleonic models

### v2

@@ IV.2 Implications for properties of nucleonic EoS
Figure 11: Posterior distributions of eight CDF parameters under astrophysical scenarios B1 and F1. The light- and dark-shaded regions represent the 68.3% and 95.4% CIs for the two-dimensional distributions, while the one-dimensional posteriors for each parameter are shown along the edges, with vertical lines indicating the 68.3% credible intervals. For comparison, results considering only the constraints from symmetric nuclear matter characteristic parameters (labeled SNM) and those including additional EFT computations (labeled SNM+EFT) are provided in the one-dimensional plots. The parameter $a_{\omega}$ (highlighted in blue) is not a free parameter in the present CDF framework.
@@
Figure 13: Posterior distributions of nine gross properties of compact stars: radius $R$ and sound speed $c_{\rm s}^{2}$ at the center of the respective 1.0, 1.4, $2.0\,M_{\odot}$ and the maximum-mass $M_{\rm max}$ configurations, under astrophysical scenarios B1 and F1. The light and dark shaded regions indicate respectively the 68.3% and 95.4% CIs of the two-dimensional distributions, while the one-dimensional posteriors for each quantity are given at the edges of the plots where vertical lines indicate the 68.3% CI credible intervals.
@@
The impact of the individual astrophysical constraints on EoS could be understood in terms of the nuclear matter characteristics at saturation density, which are shown in Fig. 9. For reference, we list here the selected characteristics predicted from the baseline scenarios B1 and F1. Scenario B1 features lower values for the incompressibility, $K_{\rm sat}=231.4_{-33.3}^{+34.4}$ MeV and negative values for the skewness, $Q_{\rm sat}=-391.0_{-170.2}^{+333.6}$ MeV (at 68.3% CI) in the isoscalar sector, and symmetry energy $J_{\rm sym}=31.24_{-0.86}^{+0.88}$ MeV and its slope $L_{\rm sym}=46.73_{-5.67}^{+5.94}$ MeV in the isovector sector. The scenario F1 has higher values for $K_{\rm sat}=244.8_{-34.7}^{+36.4}$ MeV and allows for positive value for $Q_{\rm sat}=-164.7_{-280.1}^{+548.9}$ MeV (at 68.3% CI) in the isoscalar sector, and $J_{\rm sym}=31.20_{-0.88}^{+0.89}$ MeV and $L_{\rm sym}=47.86_{-6.27}^{+6.46}$ MeV in the isovector sector. Note that in this scenario the value of $Q_{\rm sat}$ can further reach up to 900 MeV at a 95.4% CI. The correlations among nuclear characteristics at saturation density will be accessed below in Subsec. IV.3.
@@
Including additional constraints, it is seen that the lowest-order parameters $E_{\rm sat}$ and $J_{\rm sym}$ are almost identical among scenarios B or F. The parameter $K_{\rm sat}$ is reduced by 5-10 MeV ($<5\%$) and $L_{\rm sym}$ is reduced by only 1-2 MeV ($<5\%$) among scenarios B or F. These reductions are responsible for the softness required by HESS J1731 data in scenarios B2 or F2. In contrast, for scenarios B3 and F3 where the PSR J0952 data is included, such reductions mainly countervail the stiffness induced by the higher values of $Q_{\rm sat}$ and $K_{\rm sym}$ at low density regime, which now are increased up to $Q_{\rm sat}=-215.0_{-253.0}^{+435.7}$ MeV and $K_{\rm sym}=-96.8_{-16.6}^{+24.0}$ MeV (at 68.3% CI) for scenario B3, and to $Q_{\rm sat}=81.0_{-386.6}^{+548.4}$ MeV and $K_{\rm sym}=-83.8_{-19.5}^{+26.9}$ MeV for scenario F3. Note that in our modeling, the parameters $K_{\rm sat}$ and $Q_{\rm sat}$ are quasi-independent of each other, which is in contrast to previous studies of this type CDF with simplified functions (3) for density dependence of isoscalar meson-nucleon couplings [18, 24, 28]. Since the isovector parameter $K_{\rm sym}$ is not entirely independent, its variations arise from slight adjustments in the two lower-order parameters, $J_{\rm sym}$ and $L_{\rm sym}$.
@@
In the lower panels of Fig. 8 the horizontal bands indicate the nucleonic direct Urca (DU) threshold. This band is, however, model dependent [115]. For $\mu^{-}$ free case the threshold value is 11.1%; in the limit of massless $\mu$-s, which is applicable for high densities matter, it yields an upper limit of 14.8%. As seen in this plot, due to the softness (dominated by $\chi$EFT constraint) of the nucleonic EoS at low densities, the DU process will be mostly disallowed in CSs with $M\leq 2\,M_{\odot}$ if utilizing the current CDF-based EoS. The same conclusion was reached in Refs. [18, 14, 32], where the density dependence of the $\rho$-meson coupling was also modeled using an exponential form. However, note that the DU threshold may be smoothed out due to short-range correlations which allow for finite width of the baryon spectral function in dense matter [116]. Moreover, non-nucleonic DU processes may still occur if hyperons or $\Delta$-resonances appear in the inner cores of compact stars; see, for example, Refs. [94, 117, 118]. Alternatively, modifications to the density dependence of the $\rho$-meson coupling function can lead to an earlier onset of nucleonic DU processes, as discussed in Refs. [19, 23, 29, 32].
@@
In the upper panels of Fig. 10, we show the constraints for pressure obtained from heavy ion collision data. These include: (a) the constraint obtained by comparing measurements of collective flow from heavy-ion collisions with results from hadronic transport simulations using a set of EoS (labeled cFlow) [119]; (b) the constraint obtained by reproducing in hadronic transport simulations with mean-field interactions the ratios of experimentally measured kaon yields in Au+Au and C+C collisions (Kanos) [120, 121], and (c) constraint obtained from the elliptic flow data measured by the FOPI collaboration together with simulations from isospin quantum molecular dynamics (QMD) (eFlow) [122]. It is seen that the 68.3% CI region of the EoS from scenarios from B1 to B4 largely coincides with the HIC constraints, while the 95.4% CI region of the EoS from scenarios F1-F4 overlap with the constraints. Note, however, that these HIC constraints exhibit some model dependence, especially regarding the high-density behavior explored in collision simulations. Moreover, certain nuclear models used to interpret the experimental data do not support two-solar-mass CSs.
@@ IV.3 Correlations of CDF parameters, nuclear characteristics, and compact-star properties
We now focus on the correlations of CDF parameters, nuclear matter characteristics and compact-star properties, and the correlations between two of them for representative scenarios. In Fig. 11, we show the posterior distribution on the CDF parameters for scenarios B1 and F1, respectively. For each, we include the seven adjustable parameters $(g_{\sigma},\,g_{\omega},\,g_{\rho},\,a_{\sigma},\,d_{\sigma},\,d_{\omega},\,a_{\rho})$ that describe the CDF and in addition the parameter $a_{\omega}$. In these corner plots, we show the two-dimensional (2D) correlated probability distributions and along the diagonal one-dimensional (1D) probability distribution functions (PDFs). The elliptical nature of the 2D contours for a pair of parameters indicates the correlation existing between them, while a circular nature indicates no correlation. In the 1D plots, we also display the results that applied only the constraints on the SNM characteristics at saturation density (labeled SNM) as well as those which combine these with the $\chi$EFT computations (labeled SNM+EFT).
As anticipated the pair of parameters $(g_{\sigma},\,g_{\omega})$ display in Fig. 11 high correlation, as these are the parameters that determine the binding energy at saturation density. The pairs of $(d_{\sigma},\,d_{\omega})$ and $(a_{\sigma},\,a_{\omega})$ that control the density dependence of $\sigma$ and $\omega$ mesons’ couplings are also strongly correlated as they are constrained by the conditions (4). In addition, the two parameters $(g_{\rho},\,a_{\rho})$ in the isovector sector were seen to be correlated already when only the $\chi$EFT constraint is applied, and their PDFs are almost identically to those for scenarios B1 and F1, which means that the astrophysical constraints do not introduce any noticeable modifications.
@@
Now, let us examine the correlations among compact-star properties. Figure 13 presents the posterior distributions for nine key properties of CSs, including the radius $R$ and the squared sound speed $c_{\rm{s}}^{2}=dP/d\varepsilon$ at the center of stars with masses of 1.0, 1.41.01.41.0,\,1.41.0 , 1.4, and $2.0\,M_{\odot}$, as well as for the maximum-mass $M_{\rm max}$ configuration, under scenarios B1 and F1. The value of $c_{\rm{s}}^{2}$ characterizes the stiffness of the EoS at the stellar center while $R$ reflects the overall stiffness of the EoS from the core to the surface. As shown in Fig. 13, the closer the masses of two stars, the stronger the correlation between their radii or central sound speeds. This trend is particularly evident for the present CDF-based EoS models, which exhibit a relatively straightforward behavior.
@@
Correlations between nuclear characteristics and compact-star properties are encoded in the Pearson coefficients $r$. We find, notably, that the previously reported correlation between the slope of the symmetry energy $L_{\rm sym}$ and the radius of a $1.4\,M_{\odot}$ star $R_{1.4}$ [129, 130, 131, 96], appears to have largely vanished, with $r\sim 0.1$, as also observed in Refs. [132, 31]. However, a moderate correlation remains between $L_{\rm sym}$ and the radius of a $1.0\,M_{\odot}$ star - $R_{1.0}$, with a Pearson coefficient of approximately $r\sim 0.4$. It is important to note that the correlations reported in Refs. [129, 96] were observed in a set of models where the isoscalar sector was nearly fixed and only the isovector parameters varied significantly.
V Conclusions and outlook
A systematic analysis is performed for typical astrophysical scenarios where the importance of the individual impact of PSR J0952-0607 and HESS J1731-347 are elucidated by assessing the posterior distributions for key properties of CSs, nuclear characteristics at saturation density, and correlations among them, and the density dependence of EoS and the symmetry energy.
In conclusion, this work underscores the importance of combining multimessenger observations and advanced theoretical frameworks to enhance the fidelity of nuclear astrophysics models of CS. The CDF models with density-dependent meson-nucleon couplings provide a robust theoretical framework to interpret currently known gross properties of CS. Due to the limited degrees of freedom in the isovector sector of the current CDF formulation, the predicted lower limit for the radii of low-mass stars cannot be reduced much below 12 km. Consequently, refining and extending the density dependence of CDF parameters is necessary to accurately model potential ultra-compact low-mass stars within the nucleonic paradigm. (Recall that a phase transition to quark matter may be an alternative solution as indicated above, see Refs. [70, 45] and references therein). Additional parameter(s) could also be useful within the CDF framework to provide greater control over the symmetry energy’s high-density behavior, particularly through the parameter $K_{\rm sym}$. This issue will be further investigated in an upcoming study.
Acknowledgments
Appendix A Key quantities of nuclear matter and compact stars

### v3

@@ IV.2 Implications for properties of nucleonic EoS
Figure 11: Posterior distributions of eight CDF parameters under astrophysical scenarios B1 and F1. The light- and dark-shaded regions represent the 68.3% and 95.4% CIs for the two-dimensional distributions, while the one-dimensional posteriors for each parameter are shown along the edges, with vertical lines indicating the 68.3% CI. For comparison, results considering only the constraints from symmetric nuclear matter characteristic parameters (labeled SNM) and those including additional EFT computations (labeled SNM+EFT) are provided in the one-dimensional plots. The parameter $a_{\omega}$ (highlighted in blue) is not a free parameter in the present CDF framework.
@@
Figure 13: Posterior distributions of nine gross properties of compact stars: radius $R$ and sound speed $c_{\rm s}^{2}$ at the center of the respective 1.0, 1.4, $2.0\,M_{\odot}$ and the maximum-mass $M_{\rm max}$ configurations, under astrophysical scenarios B1 and F1. The light and dark shaded regions indicate respectively the 68.3% and 95.4% CIs of the two-dimensional distributions, while the one-dimensional posteriors for each quantity are given at the edges of the plots where vertical lines indicate the 68.3% CI.
@@
The impacts of the individual astrophysical constraints on EoS could be understood in terms of the nuclear matter characteristics at saturation density, which are shown in Fig. 9. For reference, we list here the selected characteristics predicted from the baseline scenarios B1 and F1. Scenario B1 features lower values for the incompressibility, $K_{\rm sat}=231.4_{-33.3}^{+34.4}$ MeV and negative values for the skewness, $Q_{\rm sat}=-391.0_{-170.2}^{+333.6}$ MeV (at 68.3% CI) in the isoscalar sector, and symmetry energy $J_{\rm sym}=31.24_{-0.86}^{+0.88}$ MeV and its slope $L_{\rm sym}=46.73_{-5.67}^{+5.94}$ MeV in the isovector sector. The scenario F1 has higher values for $K_{\rm sat}=244.8_{-34.7}^{+36.4}$ MeV and allows for positive value for $Q_{\rm sat}=-164.7_{-280.1}^{+548.9}$ MeV (at 68.3% CI) in the isoscalar sector, and $J_{\rm sym}=31.20_{-0.88}^{+0.89}$ MeV and $L_{\rm sym}=47.86_{-6.27}^{+6.46}$ MeV in the isovector sector. Note that in this scenario the value of $Q_{\rm sat}$ can further reach up to 900 MeV at a 95.4% CI. The correlations among nuclear characteristics at saturation density will be accessed below in Subsec. IV.3.
@@
Including additional constraints, it is seen that the lowest-order parameters $E_{\rm sat}$ and $J_{\rm sym}$ are almost identical among scenarios B (F). The parameter $K_{\rm sat}$ is reduced by 5-10 MeV ($<5\%$) and $L_{\rm sym}$ is reduced by only 1-2 MeV ($<5\%$) among scenarios B (F). These reductions are responsible for the softness required by HESS J1731 data in scenarios B2 or F2. In contrast, for scenarios B3 and F3 where the PSR J0952 data are included, such reductions mainly countervail the stiffness induced by the higher values of $Q_{\rm sat}$ and $K_{\rm sym}$ at low density regime, which now are increased up to $Q_{\rm sat}=-215.0_{-253.0}^{+435.7}$ MeV and $K_{\rm sym}=-96.8_{-16.6}^{+24.0}$ MeV (at 68.3% CI) for scenario B3, and to $Q_{\rm sat}=81.0_{-386.6}^{+548.4}$ MeV and $K_{\rm sym}=-83.8_{-19.5}^{+26.9}$ MeV for scenario F3. Note that in our modeling, the parameters $K_{\rm sat}$ and $Q_{\rm sat}$ are quasi-independent of each other, which is in contrast to previous studies of this type CDF with simplified functions (3) for density dependence of isoscalar meson-nucleon couplings [18, 24, 28]. Since the isovector parameter $K_{\rm sym}$ is not entirely independent, its variations arise from slight adjustments in the two lower-order parameters, $J_{\rm sym}$ and $L_{\rm sym}$.
@@
In the lower panels of Fig. 8 the horizontal bands indicate the nucleonic direct Urca (DU) threshold. This band is, however, model dependent [114]. For $\mu^{-}$ free case the threshold value is 11.1%; in the limit of massless $\mu$-s, which is applicable for high densities matter, it yields an upper limit of 14.8%. As seen in this plot, due to the softness (dominated by $\chi$EFT constraint) of the nucleonic EoS at low densities, the DU process will be mostly disallowed in CSs with $M\leq 2\,M_{\odot}$ if one utilizes the current CDF-based EoS. The same conclusion was reached in Refs. [18, 14, 32], where the density dependence of the $\rho$-meson coupling was also modeled using an exponential form. However, note that the DU threshold may be smoothed out due to short-range correlations which allow for finite width of the baryon spectral function in dense matter [115]. Moreover, non-nucleonic DU processes may still occur if hyperons or $\Delta$-resonances appear in the inner cores of compact stars; see, for example, Refs. [93, 116, 117]. Alternatively, modifications to the density dependence of the $\rho$-meson coupling function can lead to an earlier onset of nucleonic DU processes, as discussed in Refs. [19, 23, 29, 32].
@@
In the upper panels of Fig. 10, we show the constraints for pressure obtained from heavy ion collision data. These include: (a) the constraint obtained by comparing measurements of collective flow from heavy-ion collisions with results from hadronic transport simulations using a set of EoSs (labeled cFlow) [118]; (b) the constraint obtained by reproducing in hadronic transport simulations with mean-field interactions the ratios of experimentally measured kaon yields in Au+Au and C+C collisions (Kanos) [119, 120], and (c) constraint obtained from the elliptic flow data measured by the FOPI collaboration together with simulations from isospin quantum molecular dynamics (QMD) (eFlow) [121]. It is seen that the 68.3% CI regions of the EoSs from scenarios from B1-B4 largely coincide with the HIC constraints, while the 95.4% CI regions of the EoSs from scenarios F1-F4 overlap with the constraints. Note, however, that these HIC constraints exhibit some model dependence, especially regarding the high-density behavior explored in collision simulations. Moreover, certain nuclear models used to interpret the experimental data do not support two-solar-mass CSs.
@@ IV.3 Correlations of CDF parameters, nuclear characteristics, and compact-star properties
We now focus on the correlations of CDF parameters, nuclear matter characteristics and CS properties, and the correlations between two of them for representative scenarios. In Fig. 11, we show the posterior distributions on the CDF parameters for scenarios B1 and F1, respectively. For each, we include the seven adjustable parameters $(g_{\sigma},\,g_{\omega},\,g_{\rho},\,a_{\sigma},\,d_{\sigma},\,d_{\omega},\,a_{\rho})$ that describe the CDF and in addition the parameter $a_{\omega}$. In these corner plots, we show the two-dimensional (2D) correlated probability distributions and along the diagonal one-dimensional (1D) probability distribution functions (PDFs). The elliptical nature of the 2D contour for a pair of parameters indicates the correlation existing between them, while a circular nature indicates no correlation. In the 1D plots, we also display the results that applied only the constraints on the SNM characteristics at saturation density (labeled SNM) as well as those which combine these with the $\chi$EFT computations (labeled SNM+EFT).
As anticipated the pair of parameters $(g_{\sigma},\,g_{\omega})$ displayed in Fig. 11 high correlation, as these are the parameters that determine the binding energy at saturation density. The pairs of $(d_{\sigma},\,d_{\omega})$ and $(a_{\sigma},\,a_{\omega})$ that control the density dependence of $\sigma$ and $\omega$ mesons’ couplings are also strongly correlated as they are constrained by the conditions (4). In addition, the two parameters $(g_{\rho},\,a_{\rho})$ in the isovector sector were seen to be correlated already when only the $\chi$EFT constraint is applied, and their PDFs are almost identically to those for scenarios B1 and F1, which means that the astrophysical constraints do not introduce any noticeable modifications.
@@
Now, let us examine the correlations among CS properties. Figure 13 presents the posterior distributions for nine key properties of CSs, including the radius $R$ and the squared sound speed $c_{\rm{s}}^{2}=dP/d\varepsilon$ at the center of stars with masses of 1.0, 1.41.01.41.0,\,1.41.0 , 1.4, and $2.0\,M_{\odot}$, as well as for the maximum-mass $M_{\rm max}$ configuration, under scenarios B1 and F1. The value of $c_{\rm{s}}^{2}$ characterizes the stiffness of the EoS at the stellar center while $R$ reflects the overall stiffness of the EoS from the core to the surface. As shown in Fig. 13, the closer the masses of two stars, the stronger the correlation between their radii or central sound speeds. This trend is particularly evident for the present CDF-based EoS models, which exhibit a relatively straightforward behavior.
@@
Correlations between nuclear characteristics and CS properties are encoded in the Pearson coefficients $r$. We find, notably, that the previously reported correlation between the slope of the symmetry energy $L_{\rm sym}$ and the radius of a $1.4\,M_{\odot}$ star $R_{1.4}$ [128, 129, 130, 95], appears to have largely vanished, with $r\sim 0.1$, as also observed in Refs. [131, 31]. However, a moderate correlation remains between $L_{\rm sym}$ and the radius of a $1.0\,M_{\odot}$ star - $R_{1.0}$, with a Pearson coefficient of approximately $r\sim 0.4$. It is important to note that the correlations reported in Refs. [128, 95] were observed in a set of models where the isoscalar sector was nearly fixed and only the isovector parameters varied significantly.
V Conclusions and outlook
A systematic analysis is performed for typical astrophysical scenarios where the importance of the individual impacts of PSR J0952-0607 and HESS J1731-347 are elucidated by assessing the posterior distributions for key properties of CSs, nuclear characteristics at saturation density, and correlations among them, and the density dependence of EoS and the symmetry energy.
In conclusion, this work underscores the importance of combining multimessenger observations and advanced theoretical frameworks to enhance the fidelity of nuclear astrophysical models of CSs. The CDF models with density-dependent meson-nucleon couplings provide a robust theoretical framework to interpret currently known gross properties of CS. Due to the limited degrees of freedom in the isovector sector of the current CDF formulation, the predicted lower limit for the radii of low-mass stars cannot be reduced much below 12 km. Consequently, refining and extending the density dependence of CDF parameters is necessary to accurately model potential ultra-compact low-mass stars within the nucleonic paradigm. (Recall that a phase transition to quark matter may be an alternative solution as indicated above, see Refs. [70, 45] and references therein). Additional parameter(s) could also be useful within the CDF framework to provide greater control over the symmetry energy’s high-density behavior, particularly through the parameter $K_{\rm sym}$. This issue will be further investigated in an upcoming study.
Acknowledgments
Appendix A Key quantities of nuclear matter and compact stars